# A Survey of Self-Evolving Agents: On Path to Artificial Super Intelligence

### v3

@@ 5 How to Evolve
This chapter aims to systematically map and analyze the major families of self-evolution methods, providing a unified framework for understanding their principles, mechanisms, and interactions. We begin with reward-based evolution, which centers on the design of reward signals—ranging from natural language feedback and internal confidence metrics to external or implicit signals—to guide iterative self-improvement. Next, we examine imitation and demonstration learning, where agents improve by learning from high-quality exemplars, either self-generated or provided by other agents or external sources. This paradigm is particularly powerful when demonstrations are abundant or can be autonomously synthesized, and it has driven significant progress in both reasoning and multimodal domains. Finally, we introduce population-based and evolutionary methods, which draw inspiration from biological evolution and collective intelligence. These approaches maintain populations of agent variants or collaborating agents, leveraging mechanisms such as selection, mutation, crossover, and competition to explore the solution space in parallel, foster diversity, and enable the emergence of novel strategies or architectural innovations.
@@ 5.1 Reward-based Self-Evolution
Textual Feedback leverages the native modality of LLMs—natural language—to provide detailed, interpretable instructions for refinement. Unlike scalar rewards, textual feedback encapsulates nuanced critiques and actionable suggestions. Recent frameworks such as Reflexion (shinn2023reflexion, ), AdaPlanner (sun2023adaplanner, ), AgentS2 (agashe2025agents2, ), SELF (lu2023self, ), Self-Refine (madaan2023self_refine, ), SCoRe (kumar2024training, ), PAG (jiang2025pag, ), and TextGrad (yellamraju2024textgrad, ) exemplify this direction. For instance, Reflexion proposes “verbal reinforcement learning,” where agents reflect in natural language on their past trials, storing these reflections as episodic memory to guide future decisions. AdaPlanner enables closed-loop adaptive planning by allowing LLM agents to revise their plans based on both in-plan and out-of-plan feedback, while also mitigating hallucination via code-style prompts and leveraging skill discovery. Self-Refine and SELF further explore iterative self-feedback and self-correction, demonstrating that even state-of-the-art models can be improved via multi-turn, language-based self-critique, without additional supervised data or external reinforcement. Such frameworks highlight the power of language as a reward channel, enabling nuanced, flexible, and sample-efficient self-improvement.
@@
Internal Confidence-based rewards move away from external signals and instead exploit internal metrics such as the model’s probability estimates or certainty. This paradigm leverages the model’s intrinsic understanding to guide improvement without relying on external supervision. Methods such as Confidence-Informed Self-Consistency (CISC) (taubenfeld2025cisc, ), Self-Ensemble (xu2025selfensemblemitigatingconfidencedistortion, ), Self-Rewarding Self-Improving (simonds2025selfrewardingselfimproving, ), scalable best-of-N selection via self-certainty (kang2025scalablebestofnselectionlarge, ), and Self-Rewarding Language Models (yuan2025selfrewardinglanguagemodels, ) allow models to self-evaluate and calibrate their responses based on internal confidence metrics. For example, CISC weights reasoning paths by confidence scores to improve both accuracy and computational efficiency, effectively filtering high-quality solutions from multiple candidates. Self-Ensemble mitigates confidence distortion by dividing choices into smaller, more manageable groups and aggregating predictions to reduce overconfidence bias. Self-Rewarding Language Models demonstrate that models can act as their own reward function, generating training data through self-instruction and self-evaluation cycles. These approaches can reduce reliance on human labels and external evaluators, enabling scalable and autonomous self-improvement loops that can operate continuously without human intervention.
@@
External Rewards are derived from sources outside the model, such as the environment, majority voting, or explicit rules. Majority voting (shafayat2025largereasoningmodelsselftrain, ; wei2025unsupervisedposttrainingmultimodalllm, ; zhang2025consistentpathsleadtruth, ) uses consensus among multiple model outputs as a proxy for correctness, providing a self-generated but grounded reward signal. Environment feedback, including tool-based signals, is central to agentic LLM research (e.g., SWE-Dev (du2025swedevevaluatingtrainingautonomous, ), SICA (robeyns2025self, ), Feedback Friction (jiang2025feedbackfriction, ), USEagent (applis2025useagent, ), DYSTIL (wang2025dystildynamicstrategyinduction, )), where agents learn through direct interaction with real-world environments and tools. Rule-based rewards (wang2025otcpo, ; wang2025autorule, ; zhang2025learning, ; simonds2025ladder, ; wang2025ragen, ; liu2025spiral, ) use explicit constraints or logical rules as verifiable signals, particularly effective in the domains of mathematical reasoning, game play, and structured problem solving. These methods offer objective, reliable supervision but may require significant engineering or be limited in expressiveness.
@@
Implicit Reward frameworks hypothesize that LLMs can learn from feedback signals even when not explicitly labeled as rewards. For instance, “Reward Is Enough” (song2025reward, ) demonstrates that LLMs can perform in-context reinforcement learning using simple scalar signals embedded in the context window, improving their responses over rounds without explicit RL fine-tuning or supervision. This reveals an inherent capacity for models to interpret and learn from implicit feedback cues present in their input context. Recent work has expanded this concept by showing that LLMs inherently encode reward-like signals through their standard training objectives. Endogenous reward (li2025generalist, ) reveal that standard next-token prediction implicitly learns a generalist reward function, which can be extracted from model logits without additional training. Moreover, ImPlicit Self-ImprovemenT (PIT) framework (wang2024enabling, ) implicitly learns the improvement goal from human preference data without extra human efforts by maximizing the quality gap of the response conditioned on a reference response. Unlike rule-based or environment-derived external rewards, implicit reward methods offer unique advantages by discovering and utilizing reward signals that are inherently present in language modeling.
@@ 5.2 Imitation and Demonstration Learning
Imitation and demonstration learning is a paradigm in which self-evolving agents improve their capabilities by learning from high-quality exemplars, which may be generated by the agents themselves, other agents, or external sources. Unlike reward-based methods that rely on explicit reward signals, imitation-based approaches focus on reproducing and refining successful behavioral patterns through iterative self-training and bootstrapping mechanisms. This approach is particularly effective when high-quality demonstrations are available or can be autonomously generated, allowing agents to bootstrap their capabilities with minimal external supervision.
@@ 5.2.1 Self-Generated Demonstration Learning
Bootstrapping Reasoning Capabilities. zelikman2022star introduces the foundational framework for self-generated demonstration learning, enabling language models to bootstrap their reasoning capabilities through iterative self-training. This process involves generating reasoning chains for problems, fine-tuning on correct solutions, and repeating this cycle to progressively improve performance without the need for ground-truth reasoning paths. Building on this framework, recent advancements have refined the bootstrapping process through more sophisticated training strategies. For instance, hosseini2024v proposes a verifier-guided self-training approach, where separate verifier models assess the quality of generated reasoning chains before they are incorporated into the training data, enhancing the reliability of self-improvement. Additionally, koh2025adastar introduces adaptive data sampling strategies that dynamically adjust the composition of training data based on model performance across various reasoning tasks, thereby mitigating overfitting to specific problem types.
Multimodal Self-Training. Extending self-training to multimodal domains presents unique challenges in generating high-quality demonstrations that span both visual and textual modalities. deng2024enhancing demonstrates how vision-language models can improve iteratively by training on their own generated image descriptions and visual reasoning chains. The approach leverages the model’s existing visual understanding to generate detailed image descriptions, which are subsequently used to fine-tune the model’s visual perception in a bootstrapping manner. zhao2024genixer builds on this concept by empowering multimodal large language models to serve as powerful data generators, producing diverse training examples across different modalities and tasks through advanced prompt engineering and quality filtering mechanisms.
@@ 5.2.2 Cross-Agent Demonstration Learning
Multi-Agent Bootstrapped Reasoning. zhao2025sirius presents a framework for multi-agent systems to learn from each other’s successful demonstrations through bootstrapped reasoning. The system maintains an experience library containing successful interaction trajectories generated by different agents, facilitating efficient knowledge sharing and collaborative improvement. Each agent can leverage the collective experience of the entire system, thereby accelerating the learning process and enabling the discovery of diverse solution strategies. This framework illustrates how agents can specialize in different aspects of complex tasks while benefiting from the accumulated knowledge of the entire system.
Domain-Specific Demonstration Learning. Domain-specific applications of demonstration learning have proven especially effective in specialized fields where expert knowledge can be effectively transferred through demonstrations. In recommendation systems, techniques such as self-optimized fine-tuning tang2025bridging enable LLM-based recommender systems to learn from their own successful recommendation patterns, creating a feedback loop that enhances personalization over time. The system generates high-quality recommendation demonstrations from successful user interactions and uses these to fine-tune the underlying language model, ultimately leading to more accurate and personalized recommendations.
@@ 5.2.3 Hybrid Demonstration Learning
Recursive Self-Improvement. qu2024recursive demonstrates how agents can be trained to systematically improve their behavior through structured self-reflection and demonstration generation. This approach enables language model agents to introspect on their reasoning processes, identify areas for improvement, and generate corrective demonstrations to address these weaknesses. This recursive process establishes a continuous improvement loop, where agents become increasingly skilled at self-diagnosis and self-correction, leading to more robust and adaptable behavior.
Confidence-Guided Demonstration Selection. Recent developments have focused on more sophisticated mechanisms for selecting high-quality demonstrations from both self-generated and external sources. Confidence-based approaches li2024confidence utilize the model’s uncertainty estimates to determine which demonstrations are most likely to contribute positively to learning, filtering out potentially detrimental or low-quality examples. This method addresses a critical challenge in demonstration learning: poor-quality demonstrations can degrade performance. By ensuring that only high-confidence, high-quality examples are used for training, this approach helps to maintain the integrity of the learning process.
@@ 5.3 Population-based and Evolutionary Methods
Population-based and evolutionary methods represent a fundamentally different paradigm for agent evolution compared to the reward-based and imitation-based approaches discussed in previous sections. While reward-based methods typically optimize individual agents through iterative reward signals and imitation learning relies on learning from demonstrations, population-based methods draw inspiration from biological evolution and collective intelligence. These approaches maintain multiple agent variants simultaneously, allowing for parallel exploration of the solution space and the emergence of diverse capabilities through mechanisms such as selection, mutation, crossover, and competitive interaction zhang2025nature . This enables broader search coverage and the discovery of novel solutions that might be missed by gradient-based optimization. This approach is particularly valuable when the solution space is complex, multimodal, or when the optimal strategy requires fundamental architectural changes rather than parameter fine-tuning.
@@ 5.3.1 Single Agent Evolution
Learning from Evolution. This paradigm draws directly from biological evolution, maintaining populations of agent variants and applying evolutionary operators to discover improved capabilities. The Darwin Gödel Machine (DGM) zhang2025darwin exemplifies this approach through open-ended evolution of self-improving agents that maintain an archive of all historical versions, enabling branching from any past "species" rather than linear optimization. The system achieves self-referential improvement by allowing agents to directly modify their own Python codebase, with evolution driven by empirical performance on coding benchmarks and parent selection balancing performance scores with novelty rewards for diverse exploration. Complementing this code-level evolution, the Nature-Inspired Population-Based Evolution (GENOME) framework zhang2025nature directly applies genetic algorithms to language model parameter evolution, maintaining populations and using crossover, mutation, and selection operators on model weights. GENOME+ extends this with particle swarm optimization concepts, adding inheritance mechanisms and ensemble methods that demonstrate gradient-free evolutionary optimization can effectively improve model capabilities through parameter space exploration.
@@
Self-Play from Multiple Rollouts. This paradigm focuses on agents improving through iterative self-competition and rollout-based learning, where agents generate multiple trajectories and learn from their own exploration. Self-Play Fine-Tuning (SPIN) chen2024self establishes the foundation by having current models compete against previous versions, creating evolutionary pressure where only improving strategies survive without external annotations. SPC chen2025spc advances this through sophisticated adversarial co-evolution, where a "sneaky generator" learns to create deceptive errors while a "step critic" evolves to detect increasingly subtle mistakes, using automated validation to sustain improvement without human step-level annotations. STL mendes2025language demonstrates self-teaching evolution through iterative lookahead search, where value models generate training data from their own exploratory rollouts, combining numerical value learning with natural language reasoning chains to bootstrap continuous improvement. These approaches share the principle of using agents’ own generated experiences as learning signals, creating self-sustaining improvement cycles that evolve without external supervision.

### v4

@@ 5 How to Evolve
Over time, research on self-evolving agents has progressed through three major paradigms—reward-based, imitation-based, and population-based evolution—each emerging to address the limitations of the previous one. Reward-based methods first closed the feedback loop through explicit signals but suffered from brittleness and high cost. Imitation-based learning stabilized evolution by leveraging high-quality demonstrations, though sometimes at the expense of exploration. Population-based evolution then extended adaptation to collective scales, emphasizing diversity and emergent coordination. Together, these paradigms outline a coherent trajectory from individual self-improvement toward collective intelligence.
This chapter aims to systematically map and analyze the major families of self-evolution methods, providing a unified framework for understanding their principles, mechanisms, and interactions. We begin with reward-based evolution, which centers on the design of reward signals—ranging from natural language feedback and internal confidence metrics to external or implicit signals—to guide iterative self-improvement.
Next, we examine imitation and demonstration learning, where agents learn by mimicking complete, high-quality behavioral exemplars (i.e., demonstrations). While traditionally sourced from human experts, in the context of self-evolving agents, these exemplars are often generated by the agent itself or by other agents. This paradigm is particularly powerful when demonstrations are abundant or can be autonomously synthesized, and it has driven significant progress in both reasoning and multimodal domains.
Finally, we introduce population-based and evolutionary methods, which draw inspiration from biological evolution and collective intelligence. These approaches maintain populations of agent variants or collaborating agents, leveraging mechanisms such as selection, mutation, crossover, and competition to explore the solution space in parallel, foster diversity, and enable the emergence of novel strategies or architectural innovations.
@@ 5.1 Reward-based Self-Evolution
Textual Feedback leverages the native modality of LLMs—natural language—to provide detailed, interpretable instructions for refinement. Unlike scalar rewards, textual feedback encapsulates nuanced critiques and actionable suggestions. Recent frameworks such as Reflexion (Shinn et al., 2023), AdaPlanner (Sun et al., 2023), AgentS2 (Agashe et al., 2025), SELF (Lu et al., 2023), Self-Refine (Madaan et al., 2023a), SCoRe (Kumar et al., 2024), PAG (Jiang et al., 2025b), and TextGrad (Yellamraju and others, 2024) exemplify this direction. For instance, Reflexion proposes “verbal reinforcement learning,” where agents reflect in natural language on their past trials, storing these reflections as episodic memory to guide future decisions. AdaPlanner enables closed-loop adaptive planning by allowing LLM agents to revise their plans based on both in-plan and out-of-plan feedback, while also mitigating hallucination via code-style prompts and leveraging skill discovery. Self-Refine and SELF further explore iterative self-feedback and self-correction, demonstrating that even state-of-the-art models can be improved via multi-turn, language-based self-critique, without additional supervised data or external reinforcement. Such frameworks highlight the power of language as a reward channel, enabling nuanced, flexible, and sample-efficient self-improvement.
@@
Internal Confidence-based rewards move away from external signals and instead exploit internal metrics such as the model’s probability estimates or certainty. This paradigm leverages the model’s intrinsic understanding to guide improvement without relying on external supervision. Methods such as Confidence-Informed Self-Consistency (CISC) (Taubenfeld et al., 2025), Self-Ensemble (Xu et al., 2025b), Self-Rewarding Self-Improving (Simonds et al., 2025), scalable best-of-N selection via self-certainty (Kang et al., 2025), and Self-Rewarding Language Models (Yuan et al., 2025b) allow models to self-evaluate and calibrate their responses based on internal confidence metrics. For example, CISC weights reasoning paths by confidence scores to improve both accuracy and computational efficiency, effectively filtering high-quality solutions from multiple candidates. Self-Ensemble mitigates confidence distortion by dividing choices into smaller, more manageable groups and aggregating predictions to reduce overconfidence bias. Self-Rewarding Language Models demonstrate that models can act as their own reward function, generating training data through self-instruction and self-evaluation cycles. These approaches can reduce reliance on human labels and external evaluators, enabling scalable and autonomous self-improvement loops that can operate continuously without human intervention. AgentEvolver (Zhai et al., 2025) proposes a comprehensive framework to improve agent training efficiency through three synergistic mechanisms: self-questioning for autonomous task generation, self-navigating for experience-guided exploration, and self-attributing for fine-grained credit assignment. In particular, its self-attributing mechanism uses an LLM’s reasoning to retrospectively assign step-wise rewards that are dense and semantically grounded for policy optimization.
@@
External Rewards are derived from sources outside the model, such as the environment, majority voting, or explicit rules. Majority voting (Shafayat et al., 2025; Wei et al., 2025c; Zhang et al., 2025j) uses consensus among multiple model outputs as a proxy for correctness, providing a self-generated but grounded reward signal. Environment feedback, including tool-based signals, is central to agentic LLM research (e.g., SWE-Dev (Du et al., 2025), SICA (Robeyns et al., 2025a), Feedback Friction (Jiang et al., 2025a), USEagent (Applis et al., 2025), DYSTIL (Wang et al., 2025c)), where agents learn through direct interaction with real-world environments and tools. Rule-based rewards (Wang et al., 2025g; Wang and Xiong, 2025; Zhang et al., 2025a; Simonds and Yoshiyama, 2025; Wang et al., 2025p; Liu et al., 2025d) use explicit constraints or logical rules as verifiable signals, particularly effective in the domains of mathematical reasoning, game play, and structured problem solving. These methods offer objective, reliable supervision but may require significant engineering or be limited in expressiveness.
@@
Implicit Reward frameworks hypothesize that LLMs can learn from feedback signals even when not explicitly labeled as rewards. For instance, “Reward Is Enough” (Song et al., 2025) demonstrates that LLMs can perform in-context reinforcement learning using simple scalar signals embedded in the context window, improving their responses over rounds without explicit RL fine-tuning or supervision. This reveals an inherent capacity for models to interpret and learn from implicit feedback cues present in their input context. Recent work has expanded this concept by showing that LLMs inherently encode reward-like signals through their standard training objectives. Endogenous reward (Li et al., 2025e) reveal that standard next-token prediction implicitly learns a generalist reward function, which can be extracted from model logits without additional training. Moreover, ImPlicit Self-ImprovemenT (PIT) framework (Wang et al., 2024i) implicitly learns the improvement goal from human preference data without extra human efforts by maximizing the quality gap of the response conditioned on a reference response. Unlike rule-based or environment-derived external rewards, implicit reward methods offer unique advantages by discovering and utilizing reward signals that are inherently present in language modeling.
In summary, reward-based evolution provides explicit optimization and strong autonomy but remains sensitive to reward design, often trading stability and safety for adaptability and openness.
@@ 5.2 Imitation and Demonstration Learning
Imitation and demonstration learning traditionally involves an agent that learns to mimic the behavior of an expert (typically a human) from a set of demonstrations. In the context of self-evolving agents, this paradigm is adapted and generalized, which is the focus of our survey. Here, the role of the "expert" is not necessarily a fixed, external entity (e.g., human) but rather any source of high-quality demonstration. In self-evolving agents, these "expert exemplars" are typically generated by the agent itself (e.g., a past successful trajectory), by other more capable agents, or synthesized from environmental interactions.
The key distinction between imitation learning and reward-based methods lies in the nature of the feedback. Imitation learning is prescriptive and exemplar-based: the agent is provided with a complete, successful guide (e.g., a full reasoning trace) and learns to reproduce this behavior. In contrast, reward-based methods are evaluative and signal-based: the agent explores on its own and receives a scalar or textual critique, forcing it to infer the path to improvement through trial-and-error and credit assignment.
Furthermore, the evolutionary mechanism differs fundamentally from population-based methods that will be introduced later. While both imitation and reward-based learning typically focus on optimizing a single agent’s improvement through iterative refinement, population-based methods evolve a collection of agents in parallel. Their progress typically comes from selection pressure across a diverse gene pool, rather than the direct knowledge transfer from an exemplar to an individual. Therefore, imitation learning occupies a unique niche: it relies on the availability of high-quality solutions to directly guide and accelerate the evolution of an individual agent, making it exceptionally powerful when such demonstrations can be reliably and autonomously generated.
@@ 5.2.1 Self-Generated Demonstration Learning
Bootstrapping Reasoning Capabilities. Zelikman et al. (2022) introduces the foundational framework for self-generated demonstration learning, enabling language models to bootstrap their reasoning capabilities through iterative self-training. This process involves generating reasoning chains for problems, fine-tuning on correct solutions, and repeating this cycle to progressively improve performance without the need for ground-truth reasoning paths. Building on this framework, recent advancements have refined the bootstrapping process through more sophisticated training strategies. For instance, Hosseini et al. (2024) proposes a verifier-guided self-training approach, where separate verifier models assess the quality of generated reasoning chains before they are incorporated into the training data, enhancing the reliability of self-improvement. Additionally, Koh et al. (2025) introduces adaptive data sampling strategies that dynamically adjust the composition of training data based on model performance across various reasoning tasks, thereby mitigating overfitting to specific problem types. The "Explore to Evolve" paradigm (Wang et al., 2025k) extends this concept to deep research web agents by proposing an automated pipeline for generating complex, verifiable training data. The framework directs an agent to first perform proactive online exploration to gather grounded information from the live web, and then to evolve a sophisticated aggregation logic to synthesize question-answer pairs that require both information-seeking and deep reasoning.
Multimodal Self-Training. Extending self-training to multimodal domains presents unique challenges in generating high-quality demonstrations that span both visual and textual modalities. Deng et al. (2024) demonstrates how vision-language models can improve iteratively by training on their own generated image descriptions and visual reasoning chains. The approach leverages the model’s existing visual understanding to generate detailed image descriptions, which are subsequently used to fine-tune the model’s visual perception in a bootstrapping manner. Zhao et al. (2024c) builds on this concept by empowering multimodal large language models to serve as powerful data generators, producing diverse training examples across different modalities and tasks through advanced prompt engineering and quality filtering mechanisms.
@@ 5.2.2 Cross-Agent Demonstration Learning
Multi-Agent Bootstrapped Reasoning. Zhao et al. (2025b) presents a framework for multi-agent systems to learn from each other’s successful demonstrations through bootstrapped reasoning. The system maintains an experience library containing successful interaction trajectories generated by different agents, facilitating efficient knowledge sharing and collaborative improvement. Each agent can leverage the collective experience of the entire system, thereby accelerating the learning process and enabling the discovery of diverse solution strategies. This framework illustrates how agents can specialize in different aspects of complex tasks while benefiting from the accumulated knowledge of the entire system.
Domain-Specific Demonstration Learning. Domain-specific applications of demonstration learning have proven especially effective in specialized fields where expert knowledge can be effectively transferred through demonstrations. In recommendation systems, techniques such as self-optimized fine-tuning (Tang et al., 2025) enable LLM-based recommender systems to learn from their own successful recommendation patterns, creating a feedback loop that enhances personalization over time. The system generates high-quality recommendation demonstrations from successful user interactions and uses these to fine-tune the underlying language model, ultimately leading to more accurate and personalized recommendations.
@@ 5.2.3 Hybrid Demonstration Learning
Recursive Self-Improvement. Qu et al. (2024) demonstrates how agents can be trained to systematically improve their behavior through structured self-reflection and demonstration generation. This approach enables language model agents to introspect on their reasoning processes, identify areas for improvement, and generate corrective demonstrations to address these weaknesses. This recursive process establishes a continuous improvement loop, where agents become increasingly skilled at self-diagnosis and self-correction, leading to more robust and adaptable behavior.
Confidence-Guided Demonstration Selection. Recent developments have focused on more sophisticated mechanisms for selecting high-quality demonstrations from both self-generated and external sources. Confidence-based approaches (Li et al., 2024b) utilize the model’s uncertainty estimates to determine which demonstrations are most likely to contribute positively to learning, filtering out potentially detrimental or low-quality examples. This method addresses a critical challenge in demonstration learning: poor-quality demonstrations can degrade performance. By ensuring that only high-confidence, high-quality examples are used for training, this approach helps to maintain the integrity of the learning process.
@@
Overall, imitation-based evolution stabilizes learning through high-quality exemplars but often trades exploration and generalization for reliability and sample efficiency.
@@ 5.3 Population-based and Evolutionary Methods
Population-based and evolutionary methods are a paradigm with a long history in improving agent behavior that complements modern learning-based techniques. This approach, drawing inspiration from biological evolution, has deep roots in AI. The concept was formalized into a practical computational tool by John Holland, whose seminal work on the Genetic Algorithm (GA) established the core operators of selection, crossover, and mutation for refining a population of solutions (Holland, 1976). Building on this, John Koza pioneered Genetic Programming (GP), a powerful extension that directly evolves executable programs or symbolic expressions, which makes it appropriate for generating agent logic. This paradigm’s power was demonstrated by automatically synthesizing novel, human-competitive results, such as patented analog electrical circuits (Koza, 2010). This success extended into diverse domains, from evolving competitive agents for strategic games like backgammon and chess (Hauptman and Sipper, 2007; Azaria and Sipper, 2005) to discovering complex, interpretable policies in agent-based simulations, such as evolving dynamic taxation rules that outperformed static, human-designed strategies (Garuccio, 2016).
This paradigm led to landmark achievements in evolving agent systems. For example, the classic "Evolved Synthetic Creatures" co-evolved agent morphology and neural controllers in a simulated 3D world, leading to the discovery of a wide variety of novel and effective locomotion strategies (Sims, 1994). Later, the influential NEAT algorithm addressed a critical challenge by demonstrating how to evolve not just the weights but the entire topology of a neural network (Stanley and Miikkulainen, 2002). This enabled the autonomous discovery of complex agent "brains" from simple initial structures, a principle that has had an enduring impact on neuroevolution. These seminal works illustrated that evolution could construct both an agent’s physical form and its complex control systems, establishing a powerful alternative to manual design.
Building on this foundation, these methods represent a different paradigm for agent evolution compared to the reward-based and imitation-based approaches discussed in previous sections. While reward-based methods typically optimize individual agents through iterative reward signals and imitation learning relies on learning from demonstrations, population-based methods maintain multiple agent variants simultaneously. This allows for parallel exploration of the solution space and the emergence of diverse capabilities through mechanisms such as selection, mutation, crossover, and competitive interaction (Zhang et al., 2025r). By leveraging parallel search and genetic variation, these methods enable broader search coverage and the discovery of novel solutions that might be missed by gradient-based optimization. This approach is particularly valuable when the solution space is complex, multimodal, or when the optimal strategy requires fundamental architectural changes rather than parameter fine-tuning.
@@ 5.3.1 Single Agent Evolution
Learning from Evolution. This paradigm draws directly from biological evolution, maintaining populations of agent variants and applying evolutionary operators to discover improved capabilities. The Darwin Gödel Machine (DGM) (Zhang et al., 2025h) exemplifies this approach through open-ended evolution of self-improving agents that maintain an archive of all historical versions, enabling branching from any past "species" rather than linear optimization. The system achieves self-referential improvement by allowing agents to directly modify their own Python codebase, with evolution driven by empirical performance on coding benchmarks and parent selection balancing performance scores with novelty rewards for diverse exploration. Recent work has further explored using LLMs themselves to implement core evolutionary operators. For instance, LLM_GP (Hemberg et al., 2024) uses the LLM to perform mutation, crossover, and selection directly on programs represented as text, leveraging the model’s innate knowledge of code to inform the evolutionary search. Similarly, open-source frameworks like CodeEvolve (Assumpção et al., 2025) have demonstrated that evolutionary coding agents can achieve state-of-the-art results on mathematical benchmarks, sometimes outperforming proprietary systems like AlphaEvolve (Novikov et al., 2025) by using an island-based genetic algorithm and inspiration-based crossover. This principle is also explored in Self-Referential Graph HyperNetworks (Pedersen et al., 2025), where networks learn to generate their own weight mutations, allowing the rate of evolution itself to become selectable and adaptable.
Beyond evolving code and architecture, this paradigm extends to evolving the model’s parameters and internal logic. The Nature-Inspired Population-Based Evolution (GENOME) framework (Zhang et al., 2025r) directly applies genetic algorithms to language model parameter evolution, maintaining populations and using crossover, mutation, and selection operators on model weights. GENOME+ (Zhang et al., 2025r) extends this with particle swarm optimization concepts, adding inheritance mechanisms and ensemble methods that demonstrate gradient-free evolutionary optimization can effectively improve model capabilities through parameter space exploration. EvoLLM-JP (Akiba et al., 2025) takes this further by using evolutionary algorithms to optimally merge multiple foundation models into a single, specialized model with superior performance. Furthermore, some frameworks create a tightly integrated feedback loop where evolution helps model fine-tuning. SOAR (Pourcel et al., 2025), for example, alternates between an evolutionary search phase to generate candidate programs and a "hindsight learning" phase that uses all attempts (both successful and failed) to generate a rich dataset for fine-tuning the agent model, creating a virtuous cycle of self-improvement.
Self-Play. Self-play is a paradigm where agents improve through iterative interaction with versions of themselves, creating a dynamic and self-sustaining learning process. Its principles were famously demonstrated by systems like AlphaZero (Silver et al., 2017), which achieved superior performance in complex games by learning entirely without human data. The core mechanism is co-evolutionary learning: as an agent improves, its opponents (past or concurrent versions of itself) also become stronger, generating a perpetual and adaptive curriculum of increasing difficulty. This avoids the stagnation that can occur when training against a fixed environment and enables the discovery of novel, emergent strategies.
This powerful principle has been adapted for LLMs and LLM Agents, enabling them to bootstrap their capabilities from zero or minimal external data. A prominent approach involves a single model or two model instances adopting distinct, co-evolving roles. For instance, Absolute Zero (Zhao et al., 2025a) and R-Zero (Huang et al., 2025b) employ a "challenger" or "proposer" agent that generates problems at the frontier of a "solver" agent’s capabilities. A more complex multi-agent dynamic is seen in Socratic-Zero (Wang et al., 2025l), where a Solver co-evolves with a powerful Teacher that creates challenges and a Generator that distills the Teacher’s strategy for scalable curriculum creation. To address the instability of purely autonomous systems, R-Few (Yu et al., 2025b) introduces a guided approach where the challenger is grounded by a small set of human examples to prevent concept drift and diversity collapse. The system improves through a closed loop where the solver is rewarded for correctness (often verified by execution) and the challenger is rewarded for posing difficult yet solvable problems, thus driving continuous improvement without external labels. Similarly, Self-Challenging Language Model Agents (Zhou et al., 2025e) establishes a framework where an agent alternates between generating and solving complex, multi-step coding tasks, using successful trajectories to fine-tune itself. The paradigm also extends to more specialized, collaborative roles, as seen in the Sol-Ver framework (Lin et al., 2025b), where an LLM co-evolves its ability to both generate code (solver) and create corresponding unit tests (verifier). Likewise, SPELL (Yang et al., 2025e) applies this principle to long-context reasoning, with a single model cyclically adopting questioner, responder, and verifier roles to provide reliable reward signals in a domain where programmatic verification is difficult.
@@
Across these approaches, improvement is driven by self-generated learning signals derived from the agent’s own trajectories. Self-Play Fine-Tuning (SPIN) (Chen et al., 2024f) establishes a foundational approach where current models compete against previous versions, creating evolutionary pressure for improvement. SPC (Chen et al., 2025c) advances this with a more sophisticated adversarial co-evolution, featuring a "sneaky generator" that creates deceptive errors and a "step critic" that learns to detect them. STL (Mendes and Ritter, 2025) demonstrates self-teaching through iterative lookahead search, where value models generate training data from their own exploratory rollouts. Recent work, such as EvoTest (He et al., 2025b), extends these ideas by introducing a gradient-free, evolutionary framework that revises an agent’s prompt, memory, and tools between episodes.
Self-play is distinguished by a unique self-improvement mechanism: an agent learns through direct interaction with variations of itself. This typically manifests in two ways: a model competes against its own past versions to drive iterative refinement (as in SPIN), or a single model adopts distinct, interacting roles, such as a "challenger" generating novel problems for a "solver" (as in Absolute Zero). This principle of learning from dynamic interaction is different from imitation-based bootstrapping. While methods like STaR (Zelikman et al., 2022) also learn from an agent’s own outputs, they do so by filtering and training on static, successful trajectories. Self-play, in contrast, generates its learning signal from the process of the game-like interaction itself, learning from relative success even when no perfect exemplar exists. This focus on a single agent’s lineage also sets it apart from broader population-based methods: instead of evolving a large, diverse population, self-play creates a highly focused evolutionary pressure between a minimal set of policies derived from the same agent.